### v3

@@ 5.3 Evaluation on Transformer
Table 3: OOD detection results on ViT-B/16 using ImageNet-1k as ID data. All values are percentages. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [24].
@@
iNat. [38] SUN [40] Places [46] Textures [5]
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
@@
(a) CIFAR-10
(b) CIFAR-100
5.4 Evaluation on Pareto Frontier of ID accuracy and OOD Detection Performance
Some existing methods negatively affect ID accuracy; however, we have found that integrating these methods with the NAP approach can mitigate or reduce this impact, achieving a more optimal balance. The NAP approach establishes an ideal equilibrium between ID classification accuracy and OOD detection efficacy (measured in AUROC), thereby positioning it favorably on the Pareto front for superior performance. This demonstrates our method’s ability to enhance OOD detection without additional costs while maintaining the model’s classification performance, as illustrated in Figures 3(a) and  3(b) using the CIFAR benchmark [20].
Table 4: Comparison with other single OOD score on CIFAR benchmarks [20] and ImageNet-1k [6] dataset. CIFAR results are averaged across 6 different OOD tasks and ImageNet-1k results are averaged across 4 different OOD tasks.
Method CIFAR-10 CIFAR-100 ImageNet-1k
FPR95 AUROC FPR95 AUROC FPR95 AUROC
MSP [11] 48.69 92.52 80.06 74.45 73.51 79.00
Energy [24] 26.59 94.63 68.29 81.23 62.39 84.91
[…]
5.5 Ablation Studies
[…]
Table 5: Ablation results on different $w$ of NAP-E on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
Table 6: Ablation results on different $w$ of NAP-R on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
$w$ CIFAR-10 CIFAR-100 ImageNet
FPR95 AUC FPR95 AUC FPR95 AUC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 17.94 95.44 47.95 88.85 47.35 85.01
0.2 11.64 97.19 40.24 91.14 44.39 86.99
0.3 9.26 97.96 34.24 92.50 41.18 88.84
0.4 9.02 98.15 32.61 92.84 37.99 90.37
0.5 9.71 98.00 35.13 92.19 35.25 91.41
0.6 11.85 97.63 42.30 90.63 34.12 91.84
0.7 14.56 97.08 49.50 88.48 35.50 91.61
0.8 19.20 96.37 57.11 86.05 40.01 90.69
0.9 22.31 95.55 63.21 83.59 48.05 89.02
1.0 26.58 94.63 68.29 81.23 58.87 86.60
$w$ CIFAR-10 CIFAR-100 ImageNet
FPR95 AUC FPR95 AUC FPR95 AUC
0.0 26.57 92.45 54.91 85.86 49.91 83.07
0.1 20.36 94.77 49.80 87.75 48.59 84.21
0.2 14.61 96.49 44.03 89.58 46.82 85.49
0.3 10.83 97.53 36.81 91.21 44.88 86.92
0.4 9.18 98.02 29.66 92.47 42.43 88.44
0.5 9.47 98.11 25.71 93.18 39.42 89.96
0.6 10.72 97.93 26.83 93.15 36.02 91.33
0.7 12.98 97.53 32.02 92.13 33.03 92.32
0.8 16.92 96.91 43.99 89.86 31.49 92.63
0.9 21.89 96.04 57.95 86.37 34.38 91.78
1.0 29.00 94.92 69.94 82.07 49.90 88.75
Table 6: Ablation results on different $w$ of NAP-R on CIFAR [20] benchmarks and ImageNet-1k [6] dataset.
Ablation on Different Parameter $w$. In Table 6 and 6, we show the various OOD detection performances of NAP-E and NAP-R on various datasets by varying the threshold $w$. It is important to note that our approach is essentially an ensemble approach. Generally, in ensemble approaches, it is advisable to avoid setting $w$ extremely close to 0 or 1, to ensure effective contribution from all constituent models especially when the models are carefully chosen, diverse and useful. As an ensemble method, we provide the results of $w$ over the entire interval [0,1] in Table 6 to show that our ensemble can always achieve significant improvements over the existing baseline and NAP. Note that these results are not provdied to not to demonstrate the sensitivity of NAP to $w$. Furthermore, Table 6 show that near-optimal (if not optimal) performance is attainable around the mid-range values, especially near 0.5. As $w$ is the only hyper parameter involved in our method, it allows us to employ a binary search to to efficiently determine a desired value for $w$. The significant improvements in ensemble results verify that our method is orthogonal to existing energy-based approaches, which is one of the main contributions of this paper to the community. See the Appendix for more details about the performance impact of changes in $w$.
@@ 5.5 Ablation Studies
Does NAP Work Across All Network Layers? To assess the efficacy of the NAP across various layers within the DenseNet architecture, we conduct a detailed examination of activation distributions utilizing the CIFAR-10 [20] and Places365 datasets [46]. Analysis focused on four pivotal points within the network: post-first convolutional layer, pre-pooling in the first and second transition blocks, and immediately preceding the final global pooling layer. Selected instances from the empirical findings are illustrated in Selected examples from our tests are shown in Figure 5, with more detailed visuals in the Appendix. From these figures, it’s clear that the deeper the layer, the easier it is to tell apart ID and OOD samples based on the maximum and average activation values. In the network’s early layers, neurons pick up basic features common to both ID and OOD samples, which explains why the lines cross in Figures 4(a) and 4(b). Moving to deeper layers, neurons start capturing more complex, meaningful features. As seen in Figures 4(c) and 4(d), ID samples with specific meanings trigger larger responses. This indicates that NAP works best closer to the end of the network, especially before the global pooling layer, highlighting its value in making models more reliable for OOD detection.

### v4

@@ 5.3 Evaluation on Transformer
Table 3: OOD detection results on ViT-B/16 using ImageNet-1k as ID data. All values are percentages.
@@
iNaturalist [45] SUN [47] Places [53] Textures [5]
FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$
@@
Discussion. We compare only with MSP [14] and Energy [27] because methods such as ASH [8] and ReAct [38] are specifically designed for CNNs and fail completely when applied to ViT. For fairness, we do not include these methods in our comparison.
5.4 Does NAP work across all network layers?
To assess the efficacy of the NAP across various layers within the DenseNet architecture, we conduct a detailed examination of activation distributions utilizing the CIFAR-10 [23] and Places365 datasets [53]. Analysis focused on four pivotal points within the network: post-first convolutional layer, pre-pooling in the first and second transition blocks, and immediately preceding the final global pooling layer. Selected instances from the empirical findings are illustrated in Selected examples from our tests are shown in Figure 4, with more detailed visuals in the Appendix C. From these figures, it’s clear that the deeper the layer, the easier it is to tell apart ID and OOD samples based on the maximum and average activation values. In the network’s early layers, neurons pick up basic features common to both ID and OOD samples, which explains why the lines cross in Figures 4(a) and 4(b). Moving to deeper layers, neurons start capturing more complex, meaningful features. As seen in Figures 4(c) and 4(d), ID samples with specific meanings trigger larger responses. This indicates that NAP works best closer to the end of the network, especially before the global pooling layer, highlighting its value in making models more reliable for OOD detection.